OpenStudy (anonymous):

Use the distributive property to factor the expression below. 54x^2y - 12x^2y^2

sam (.sam.):

factor out the GCF of 6x^(2)y from each term in the polynomial 6x^(2)y(9)+6x^(2)y(-2y) Factor out the GCF of 6x^(2)y 6x^(2)y(9-2y)

OpenStudy (anonymous):

54x^2y = 2*3*3*3*x*x*y 12x^2y^2 = 2*2*3*x*x*y*y GCF = 2*3*x*x*y = 6x^2 so we get 6x^2y (9 - 2y)